• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Deur A.

    Log entry time 16:03:11 on October 27,1998

    Entry number 3850

    keyword=run summary

    8:10 Target ion chamber total dose:179 rad.
    8:40 End |_ runs @ E=3.4 GeV, Ep=2.580 GeV, Theta=15.5 degree
    Tot # of events 27.8 M
    8:43 Compton polarimeter checked the charge asymmetry prior to change the pockel cell voltage : Aq=680 ppm.
    took also a short run @ E=3.4 GeV, Ep=2.580 GeV, Theta=15.5 degree
    8:55 change momentum to P=2.382 GeV/c. Problem with hadron arm magnets
    lost communication with hadron arm magnet. Javier came, work and find the problem (bug in the code). He will fix the problem during accelerator check
    9:00 NMR measurement: target polarisation (transverse): 33 %
    9:25 Compton checked charge asymmetry after to have change the Pockel cell current: -37 +/-24 ppm
    9:30 BCM linearity tests + BCM calibration. Ended at 11:50
    12:00 NMR: target polarisation (longitudinal) 34 %
    12:15 raster check 3*4 mm
    12:20 begin \\ runs @ E=3.4 GeV, P=2.382 GeV/c
    13:00 notice a light leak in the scintillators. Should not affect the data quality
    13:00 pedestal run
    15:15 back to // runs.